OpenStudy (anonymous):

consider the following quadratic equation: 10x^2+19x+9=3 solve the quadratic equation above by factoring and enter solution

OpenStudy (anonymous):

(2x+3)(5x+2) was correct answer from step 1

myininaya (myininaya):

10x^2+19x+9-3=0 10x^2+19x+6=0 ----------------------- lets think we need to find factors of a*c that add up to be b a*c=10*6=5*12=15*4 15+4=19 so we have 10x^2+15x+4x+6=0 5x(2x+3)+2(2x+3)=0 (5x+2)(2x+3)=0 so we have x=-2/5 or x=-3/2